In parallelogram ABCD if angle A = 5x+30, and m angle D = x, find m angle C.: In parallelogram ABCD if angle A = 5x+30, and m angle D = x, find m angle C.
Answer by Fombitz(1799)

The interior angles of the parallelogram equal 360 degrees.
Also, the oppposite angles are equal.
Angle A = Angle C
Angle B = Angle D
A + B + C + D = 360 The sum of the angles is 360.
A+D+A+D=360 Substitute for B and C.
2A+2D=360 Simplify.
A+D=180Divide by 2.
5x+30+x=180Substitute for A and D.
6x=150Simplify.
x=25Solve for x.
5x+30=155Solve for A and C.
Angle A = Angle C = 155 degrees
Angle B = Angle D = 25 degrees